Operations Report – February 18th

Crew 275 Operations Report 18Feb2023
SOL: 6

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Science Dome: 3 leaks in the walls – same as yesterday
• Propane Generator

Notes on non-nominal systems:
• Concerning toilet, see below in the last section.

Spirit rover used: No
• Hours: 209.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.1
• Beginning charge: 100%
• Ending charge: 93%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 220.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.5
• Beginning charge: 100%
• Ending charge: 81%
• Currently charging: Yes

General notes on rovers: Rovers were used for today’s EVA. Some snow was removed from the seats of Perseverance.

Summary of Hab operations: The Crew prepared the EVA in the morning and performed routine operations and experiments. The Lower Deck heater was used all day. Temperature in the Lower Deck was at 72°F all day, and the temperature in the Lower Deck was comfortable
• WATER USE: 12.87 gallons
• Water: 383.67 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: Spacesuits 1, 2 and 3 were used today, nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations, and relocated two tomato crops which were on the upper storage.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: No

Summary of ScienceDome operations: The Commander and Crew Scientist tested some experimental protocols. The Crew Scientist used a microscope to test a protocol concerning rock samples.
• Dual split: Switched off during the night according to Mission Support

Summary of RAM operations: No operations were conducted in the RAM. We retrieved the 6 AAA batteries and the user manual of the oven.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• Greenhab radio: the volume of the channel 1 radio was very low. I conducted a check on this radio with the user manual, and the volume is still low. We decided to change radios between channel 1 and 22 in the Greenhab. Indeed, important communication among the Crew are transmitted on channel 1, and transmissions with Mission Support are always managed in the Hab and summarized on channel 1.
• Concerning the toilet, I have listed this system as non-nominal since the beginning of the mission because it was reported as so before. However, we have not experienced and bad smell since the beginning of the mission. Do you still need to change the toilet tank?
• We did not experience any failure for the propane generator, is it still considered as non-nominal?

Copyright © The Mars Society. All rights reserved. | Main Site